OpenStudy (anonymous):

-8=z/3 (z over 3 ) z=8/3 z=-5 z=-24 z=3/8

OpenStudy (solomonzelman):

C

OpenStudy (anonymous):

do you know -9=5/17n n=153/5 n=5/153 n=17/45 n=45/17

OpenStudy (solomonzelman):

None of them, it is negative! it is the same as A, but negative.

OpenStudy (anonymous):

all of them are negative im sorry lol

OpenStudy (anonymous):

i forgot to put the signs

OpenStudy (solomonzelman):

So then (still) A.

OpenStudy (anonymous):

okay what about which equation is an identity ? 8-(5x+2)=-5x-6 7z+10-z=8z-2(z-5) 8m-4=5m+8-m 6y+5=6y-5

OpenStudy (solomonzelman):

A